OpenStudy (anonymous):

help Medals Which mathematical sentence represents the solution for d in the equation 6e = ef + 3d?

OpenStudy (bibby):

solve for d. start by isolating d. How would you get d (3d) alone one one side of the equation

OpenStudy (anonymous):

idk where to start

OpenStudy (bibby):

subtract ef from both sides. where does that leave you?

OpenStudy (anonymous):

6=+3d

OpenStudy (bibby):

6e = ef + 3d? -ef -ef -------------- 6e - ef = 3d

OpenStudy (anonymous):

is that it

OpenStudy (bibby):

you're solving for d, not 3d

OpenStudy (anonymous):

yes

OpenStudy (bibby):

... how do you make 3d become d?

OpenStudy (anonymous):

sorry i dont help black people

OpenStudy (anonymous):

so subtract the 3and 6

OpenStudy (bibby):

rofl

OpenStudy (bibby):

6e - ef = 3d That's 3 TIMES d. You can't subtract 3.

OpenStudy (isaiah.feynman):